I need help with this question.
Back to Assignment Attempts: Average: 74 Attention: Due to a bug in Google Chrome, this page may not function correctly. Click here to learn more. Aa Aa 15. Mortgage payments Mortgages, loans taken to purchase a property, involve regular payments at fixed intervals and are treated as reverse annuities. Mortgages are the reverse of annuities, because you get a lump-sum amount as a loan in the beginning, and then you make monthly payments to the lender Youve decided to buy a house that is valued at $1 million. You have $250,000 to use as a down payment on the house, and want to take out a mortgage for the remainder of the purchase price. Your bank has approved your $750,000 mortgage, and is offering a standard 30-year mortgage at a 9% fixed nominal interest rate (called the loans annual percentage rate or APR). Under this loan proposal, your mortgage payment will be month. (Note: Round the final value of any interest rate used to four decimal places.) Your friends ot of money on interest. If your bank approves a 15-year, $750,000 loan at a fixed nominal interest rate of 9% (APR), then the difference in the monthly payment of the 15-year mortgage and 30-year mortgage will be suggest that you take a 15-year mortgage, because a 30-year mortgage is too long and you will pay a ? (Note: Round the final value of any interest rate used to four decimal places.) It is likely that you wont like the prospect of paying more money each month, but if you do take out a 15-year mortgage, you will make far fewer payments and will pay a lot less in interest. How much more total interest will you pay over the life of the loan it you take out a 30-year mortgage Instead of a 1s-year mortgage? O $1,028,123.14 O $803,221.20 O $1,108,445.26 O $947,801.02 Which of the following statenents is not true about mortgages? O Mortgages are examples of amortized loans. O Every payment made toward an amortized loan consists of two parts-interest and repayment of principal, O If the payment is less than the interest due, the ending balance of the loan will decrease. O The ending balance of an amortized loan contract will be zero. O Type here to search
Answer #1

1-

Monthly payment on Mortgage

Using PMT function in M S excel

pmt(rate,nper,pv,fv,type) = rate = 9/12 =.75% nper =30*12 =360 pv =750000 fv =0 type =0

2-

Monthly payment on Mortgage

Using PMT function in M S excel

pmt(rate,nper,pv,fv,type) = rate = 9/12 =.75% nper =15*12 =180 pv =750000 fv =0 type =0

Difference in monthly payment

7606.9994-6034.6696

1572.3298

3-

total interest on 30 years loan

(6034.6696*360)-750000

1422481.06

total interest on 15 years loan

(7606.9994*180)-750000

619259.892

difference in interest paid

803221.164

4-

3rd is false